# Performances of a radial TPC for the detection of neutrinoless double beta decay

### v1

@@ Abstract
To search for $\beta\beta 0\nu$ decay with unprecedented sensitivity, the R2D2 collaboration is developing a radial time projection chamber with a fiducial mass of half a ton of 136Xe at high pressure. The various approaches implemented to eliminate the radioactive background are presented in terms of detector design, topological recognition of interactions, and event energy reconstruction. The developed tools allow for the disentangling of the sought-after signal from the background, and the projected sensitivity after ten years of data taking yields a half-life limit exceeding $10^{27}$ years, along with a constraint on the effective neutrino mass $m_{\beta\beta}$, sufficient to exclude the inverted mass hierarchy region.
@@ 1 Introduction
The search for neutrinoless double beta decay ($\beta\beta 0\nu$) is essential for determining whether neutrinos are Dirac or Majorana particles, i.e. whether they are distinct from their antiparticles or identical to them. Identifying neutrinos as Majorana particles would imply lepton number violation, a critical feature in the leptogenesis theory, which explains the matter-antimatter asymmetry in the universe Bodeker:2020ghk ; Agostini:2022zub . $\beta\beta 0\nu$ is the only process that can confirm the Majorana nature of neutrinos, as it occurs only if the neutrino is its own antiparticle DellOro:2016tmg . Additionally, detecting $\beta\beta 0\nu$ would provide insights into the absolute neutrino mass, an unresolved property of these particles. Thus, the search for this rare decay is pivotal for advancing both particle physics and cosmology.
@@
In parallel, the issue of low background was addressed by minimizing the total material budget of the detector, a key concept of the proposed technology, which utilizes only one readout channel. A significant advancement was achieved through the use of a thin composite vessel to contain the pressurized gas. Ongoing studies on different carbon fibers and glues used to construct this gas tank are being conducted in collaboration with industry partners, (IRT Jules Verne IRT ).
The cylindrical active volume, which is 2.5 m long with a diameter of 1 m, can accommodate a usable xenon mass of 582 kg at 40 bars. Enclosed in a tank, this simple detection system could therefore be competitive with current ton-scale projects. The cost of the proposed detector should be much lower than that of currently identified experiments and would not require cryogenic infrastructure. Preliminary sensitivity studies on a spherical TPC were published in 2018 Meregaglia:2017nhx , demonstrating that, in principle, a background-free detector at the ton scale was feasible.
In the present work, a comprehensive sensitivity study of a ton-scale CPC is presented, based on the R&D results obtained over the last six years. A sensitivity on the half-life $T^{0\nu}_{1/2}$ greater than $10^{27}$ years can be achieved after ten years of data taking, corresponding to an effective neutrino mass $m_{\beta\beta}$ in the range of 5 to 21 meV.
@@ 2 Detector Setup
The detector geometry selected for the current sensitivity study is shown in the schematic view presented in Fig. 1. The main components are described below. The mechanical structure inherently minimizes the material budget, ensuring the lowest possible internal radioactivity. Although some elements, such as gas recirculation and purification, are not directly related to the detector simulation and do not affect the sensitivity study, all components are addressed for completeness.
@@
    The detector operates in ionization mode (i.e. without avalanche at the anode). For the current study, a copper rod with a radius of 1 cm was chosen as the anode to model the collection of drifted electrons. However, in practice, a resistive anode will be used to exploit the charge sharing necessary for localization along the anode axis. The choice of radius is based on tests conducted on a previous CPC prototype, which demonstrated that a thick anode tube creates a higher drift field away from the anode, mitigating charge loss effects, and reduces geometrical distortions due to anode inhomogeneities. Improved energy resolution was indeed achieved with thick anodes in ionization mode Bouet:2024njk ; FabriceTPC . The anode is grounded and connected to the readout electronics at both ends. Grounding the anode and applying a negative high voltage to the cathode decouples the signal readout from the high voltage, thereby reducing electronic noise on the signal.
@@
    The xenon vessel design is based on the novelty of using composite materials. These materials consist of a mix of carbon fibers and epoxy glue. Composite tanks, few millimeter thick and withstanding pressures of several hundred bars, are already used today in space industry for rocket fuel tanks and in the automotive industry for hydrogen reservoirs. The structure used in the present sensitivity study is made of 1.5 cm of carbon fiber-based composite material. This is very similar to the tank currently under test, produced by MAHYTEC company MAHYTEC .
    Dedicated studies are ongoing in cooperation with IRT Jules Verne company IRT to reduce the thickness and select low radioactivity materials. Several measurements of epoxy glues for low-background experiments have already been conducted Busto:2002hq , but no one has ever used carbon fibers in low radioactivity experiments. Carbon fibers are, in principle, relatively pure and radioactivity at the level of few $\mu$Bq per kg can be achieved. However, they can become contaminated with uranium and thorium during the industrial sizing process, which prepares them for integration with the glue. Current measurements of carbon fibers and resins are being conducted in partnership with industry, aiming for a composite material with a radioactivity level of 10 $\mu$Bq/kg.
@@
To install the detector in an underground laboratory, a footprint of approximately 50 m2 is required to ensure safe operation of the detector and the recirculation/recuperation system. The estimated cost of the full detector, excluding the cost of xenon gas, is projected to be well below 5 M€.
@@ 3 Signal waveform construction
Figure 2: GARFIELD++ simulation of the drift time as a function of the radial position for a cathode of 50 cm at -20 kV and a central grounded anode with a radius of 1 cm. The CPC is filled with xenon gas at 40 bar.
@@
    Figure 3: Example waveform of a $\beta\beta 0\nu$ events starting at a radial position of 26.5 cm and a longitudinal position of Z=139 cm.
@@ 4 Waveform analysis
Figure 4: Example of the decomposition of the total signal $S$ from a single event of 208Tl produced by the 232Th decay chain simulation.. The full waveform is shown in Fig. 4(a), while a zoomed-in view is presented in Fig. 4(b).The blue line indicates the original waveform, the green line shows the decomposition result, and the cyan curve displays the residue of the decomposition. The red stars mark the maxima of the different elementary contributions identified by the decomposition algorithm.
@@
The analysis of the internal modulation of the signals can be performed to recover the radial and longitudinal location of the energy deposits, and thus reconstruct the particle track inside the gas. By analysing the different peaks, it is possible to reconstruct the radial position $R$ via the drift time and the longitudinal position $Z$ via the amplitude ratio of $Sl$ and $Sr$. Furthermore, a finer decomposition of the total signal $S$ into elementary contributions could be considered. This was carried out (see Fig. 4) by a simple iterative method of signal synthesis, where the process is interrupted when $Q_{synthesized}/Q_{r}>0.99$. As a result, it was possible to create a more detailed map of the interaction footprint in the $(R,Z)$ plane (see Fig. 5), with an accurate estimate of the initial primary charge for each GEANT4 hit being also a by-product of the routine.
@@
  • The number of peaks in the waveform $N_{p}$. As explained before, for each peak, the radial position is reconstructed based on its drift time, and its $Z$ position is determined from the ratio between the left and right waveforms.
@@ 5 Energy reconstruction
  • The synthesis method also allows calculation of the reconstructed energy deficit. For each cluster of an event, identified by the decomposition algorithm at a given radial position $R$ (i.e. at a given drift time), the energy deficit is determined by multiplying the value of the lost charge function at that position by the number of primaries estimated at the same point. Summing over the different clusters enables the correction of the energy deficit due to the varying initial radial positions of the primaries in the event (see Fig. 8).
Figure 8: Instantaneous current (blue) induced by an elementary charge and corresponding lost charge (orange) as a function of the drift time. For illustration, the marker indicates the energy lost associated with a primary created at $R$ = 35.38 cm, and the corresponding the waveform is shown by the blue area.
Figure 9: Reconstructed charge $Q_{synthesized}$ of simulated $\beta\beta 0\nu$ events using the synthesis method as a function of the signal mean radial position $R$ before and after the calibration procedure explained in the text.
@@
Given the dependence of the signal on the radial position $R$, an experimental energy calibration of the detector must be conducted. Two possible solutions are currently being studied. The first option involves using a dedicated point-like alpha source deployed at specific radial positions, though this could potentially disturb the drift electric field. Another attractive solution is to exploit the diffuse radon dissolved in the gas, as it likely constitutes an ultimate and irreducible background.
The energies reconstructed before and after the calibration procedures previously discussed are shown in Fig. 10. A Gaussian fit of the distribution obtained after energy correction shows a resolution of 0.1% FWHM for both methods. Although both approaches yield similar results, the synthesis routine is deemed more satisfactory from a methodological perspective. Consequently, this method was used for energy estimation in this article.
@@
To perform a realistic sensitivity study, the energy resolution must account for additionnal uncertainties. At this stage, effects due to fluctuations in the number of electrons or charge loss from attachment to gas impurities during drift have not been included. A separate statistical error of 0.3% FWHM is expected, assuming a Fano factor of approximately 0.15 Nygren:2011zz . Additionally, for real data, the influence of front-end electronics (amplifier response function) or digital processing (time shift of the sampling comb between the two ADCs) can also affect the energy resolution.
@@
Figure 11: Selection cuts, indicated by the vertical red dashed line, on several observables, namely: number of reconstructed peaks 11(a), $\Delta R$ 11(b), $\Delta Z$ 11(c), $\Delta R\times\Delta Z$ 11(d), and larger radial energy deposit position 11(e). The distribution of such observables (after ROI selection) for $\beta\beta 0\nu$ (blue), Thorium 232 chain (orange) and Uranium 238 chain (green) are shown.
@@ 8.2 Composite material radioactivity
A specific composite material selection process is currently underway in collaboration with the IRT Jules Verne company. Various carbon fibres are being evaluated based on their radioactivity and mechanical strength to withstand pressure, alongside different epoxy glues to bind them. As mentioned in Sec. 2, a structure with a 1.5cm thick carbon fibre-based composite material has been assumed. The decay chains of 238U and 232Th, as well as 60Co decays, were simulated uniformly within the carbon fibre layer. For the sensitivity studies, it was assumed that material radioactivity at the level of 10 $\mu$Bq/kg can be achieved. While this remains a critical point yet to be fully demonstrated, promising results have been obtained through the partnership with IRT Jules Verne. In any case, the presented background event rates per year can be scaled linearly with respect to different material radioactivity levels.
In the GEANT4 simulation, a conservative time window of 10ms was assumed to separate events in the decay chain, considering a maximal drift time of approximately 5ms. Given the half-lives of 164 $\mu$s for 214Po (from the 238U chain) and 300 ns for 212Po (from the 232Th chain), Bi-Po decays are treated as a single event in the detector simulation. For each decay chain, a sample of 10 million events was simulated. Only events depositing more than 2.2 MeV in the xenon active volume at the GEANT4 level were further processed. Based on the assumed activity of 10 $\mu$Bq/kg and the composite mass of approximately 330 kg, the generated statistics correspond to about 10 years of data for 232Th, 7 years for 238U, and 90 years for 60Co.
The total number of expected events per year in the ROI, before any selection cuts, can be computed, as well as the irreducible background events surviving the selection cuts described in Sec.6. A summary is presented in Tab. 2, which shows that the total number of background events from the composite vessel’s radioactivity is approximately 1.3 per year. All the selected events in the 232Th decay chain originate from 208Tl decays, while the events from the 238U decay chain are from 214Bi decays. In both cases, a $\gamma$ deposits all the energy in a confined region of the detector, with no multi-Compton topology to enable event rejection.

### v2

# Simulation of a radial TPC for the detection of neutrinoless double beta decay

@@ Abstract
To search for $\beta\beta 0\nu$ decay with unprecedented sensitivity, the R2D2 collaboration is developing a radial time projection chamber with a fiducial mass of half a tonne of 136Xe at high pressure. The various approaches implemented to eliminate the radioactive background are presented in terms of detector design, topological recognition of interactions, and event energy reconstruction. The developed tools enable the disentangling of the sought-after signal from the background. The projected sensitivity after ten years of data taking yields a half-life limit exceeding $10^{27}$ years, along with a constraint on the effective neutrino mass $m_{\beta\beta}$ that could cover a large fraction of the inverted mass hierarchy region, depending on the final experimental background.
@@ 1 Introduction
The search for neutrinoless double beta decay ($\beta\beta 0\nu$) is essential for determining whether neutrinos are Dirac or Majorana particles, i.e. whether they are distinct from their antiparticles or identical to them. Identifying neutrinos as Majorana particles would imply lepton number violation, a critical feature of leptogenesis models that explain the matter-antimatter asymmetry in the universe Bodeker:2020ghk ; Agostini:2022zub . $\beta\beta 0\nu$ is the only process that can confirm the Majorana nature of neutrinos, as it occurs only if the neutrino is its own antiparticle DellOro:2016tmg . Additionally, detecting $\beta\beta 0\nu$ would provide insights into the absolute neutrino mass, an unresolved property of these particles. Thus, the search for this rare decay is pivotal for advancing both particle physics and cosmology.
@@
In parallel, the issue of low background was addressed by reducing the total material budget of the detector, a central consideration in the design process, which utilizes a single readout channel. A significant advancement could be achieved through the use of a thin composite vessel to contain the pressurized gas. Ongoing studies on different carbon fibers and glues used to construct this gas tank are being conducted in collaboration with industry partners, (IRT Jules Verne IRT ).
The cylindrical active volume, which is 2.5 m long with a diameter of 1 m, can accommodate an active xenon mass of 582 kg at 40 bars. Enclosed in a tank, this simple detection system could therefore be competitive with current ton-scale projects. The cost of the proposed detector should be much lower than that of currently identified experiments and would not require cryogenic infrastructure. Preliminary sensitivity studies on a spherical TPC were published in 2018 Meregaglia:2017nhx , demonstrating that, in principle, a background-free detector at the ton scale was feasible.
In the present work, a comprehensive sensitivity study of a ton-scale CPC is presented, based on the R&D results obtained over the last six years. A sensitivity on the half-life $T^{0\nu}_{1/2}$ greater than $10^{27}$ years can be achieved after ten years of data taking, corresponding to an effective neutrino mass $m_{\beta\beta}$ in the range of 13 to 57 meV. Such a limit can be reduced to 8–35 meV if the ultimate R2D2 goal of reaching zero background is achieved.
@@ 2 Detector Setup
The detector geometry selected for the current sensitivity study is shown in the schematic view presented in Fig. 1. The main components are described below. The mechanical structure inherently minimizes the material budget, ensuring the lowest possible instrumental radioactivity. Although some elements, such as gas recirculation and purification, are not directly related to the detector simulation, all components are addressed for completeness..
@@
    The detector operates in ionization mode (i.e. without avalanche at the anode). A detailed study on the anode is ongoing to minimise its radioactivity. The most promising option, currently under testing and considered in the following simulation, is a hollow tube with a thickness of 1 mm and an external radius of 1 cm, made of polymer materials coated with a resistive foil a few micrometres thick. The choice of radius is based on tests conducted on a previous CPC prototype, which demonstrated that a thick anode tube creates a higher drift field away from the anode, mitigating charge loss effects, and reduces geometrical distortions due to anode inhomogeneities. Improved energy resolution was indeed achieved with thick anodes in ionization mode Bouet:2024njk ; FabriceTPC . The anode is grounded and connected to the readout electronics at both ends. Grounding the anode and applying a negative high voltage to the cathode decouples the signal readout from the high voltage, thereby reducing electronic noise on the signal.
@@
    The xenon vessel design is based on the use of composite materials. These materials consist of a mix of carbon fibers and epoxy glue. Composite tanks, few millimeter thick and withstanding pressures of several hundred bars, are already used today in space industry for rocket fuel tanks and in the automotive industry for hydrogen reservoirs. The structure used in the present sensitivity study is made of 1.5 cm of carbon fiber-based composite material. This is very similar to the tank currently under test, produced by MAHYTEC company MAHYTEC .
    Dedicated studies are ongoing in cooperation with IRT Jules Verne company IRT to reduce the thickness and select low radioactivity materials. Several measurements of epoxy glues for low-background experiments have already been conducted Busto:2002hq , and some measurements of carbon fibres were also carried out, for example, by the nEXO collaboration nEXO:2018ylp . While carbon fibres are inherently low in contaminants, they can become contaminated with uranium and thorium during the industrial sizing process, which prepares them for integration with the glue. Current measurements of carbon fibers and resins are being conducted in partnership with industry, aiming for a composite material with a radioactivity level of 10 $\mu$Bq/kg.
@@
To install the detector in an underground laboratory, a footprint of approximately 50 m2 is required to ensure safe operation of the detector and the recirculation/recuperation system.
@@ 3 Signal waveform construction
Figure 2: GARFIELD++ simulation of the drift time as a function of the radial position for a cathode of 50 cm at -20 kV and a central grounded anode with a radius of 1 cm. The CPC is filled with xenon gas at 40 bars.
Figure 3: Example waveform of a $\beta\beta 0\nu$ events starting at a radial position of 26.5 cm and a longitudinal position of Z=139 cm. The start (not shown) of the signal begins at bin 0. Each time sample corresponds to 0.5 $\mu$s given the sampling at 2 MHz.
@@ 4 Waveform analysis
Figure 4: Example of the decomposition of the total signal $S$ of the same $\beta\beta 0\nu$ simulated event shown in Fig. 3. The result after the first iteration is shown in Fig. 4(a), whereas the result at the end of the process is presented in Fig. 4(b).The blue line indicates the original waveform, the green line shows the decomposition result, and the cyan curve displays the residue of the decomposition. The red stars mark the original amplitude of the different elementary contributions identified by the decomposition algorithm, whereas the black dots represent the amplitude of the contributions after the decomposition.
@@
The analysis of the internal modulation of the signals can be performed to recover the radial and longitudinal location of the energy deposits, and thus reconstruct the particle track inside the gas. Due to the spiked shape of the induced current in the radial TPC (see Fig. 10 in Sec. 5 and Ref. Bouet:2024njk ), a local maximum in the signal is generated each time a primary reaches the anode (or, within the limits of the time resolution, a cluster of primaries). In this approach, the amplitude of each peak is directly related to the amount of charge reaching the anode. Regarding radial localization, the sampling index of the peak, subtracted from that of the signal’s start (common to all peaks in the same event), determines the drift time associated with the peak (calculated by multiplying by the sampling period). The correspondence between this time lapse and the radial position of the energy deposition for the primary is provided in Fig. 2, enabling the extraction of a radial position $R$ for the peak. Consequently, referencing an external time stamp is no longer necessary. The longitudinal position $Z$ of a peak is determined by the principle of charge sharing, using the ratio of the left ($S_{l}$) and right ($S_{r}$) amplitudes of the peak, as described in Sec. 3.
Building upon previous calculations, a complete decomposition of the total signal $S$ into elementary contributions of induced current has been developed. This process is achieved (see Fig. 4) using a simple iterative method for signal decomposition and synthesis. The decomposition involves identifying peaks in the original waveform and scaling the elementary current template (truncated if necessary) in amplitude to match each peak. The synthesized waveform is then formed by summing these scaled contributions. Any remaining peaks not previously identified are searched for in the residual waveform resulting from the subtraction of the synthesized waveform from the original one. The process continues iteratively until the ratio of the synthesized charge to the original charge, $Q_{synthesized}/Q_{r}$, exceeds 0.99.
Figure 5: Example of a signal decomposition (in full time range) showing multiple distant interactions, originating from a single 208Tl event produced by the simulation of the 232Th decay chain.
The main advantage of this method is that it allows for a more accurate and detailed map of the interaction footprint in the $(R,Z)$ plane (see Fig. 5 and 6). Furthermore, the decomposition allows for extracting the true amplitude of each peak, making it possible to estimate the actual number of primaries associated with each. Finally, through synthesis, the unobserved portion of the signal, which depends on the initial radial position of each primary, becomes quantifiable. This enables the reconstruction of the true energy of the signal, determined by integrating both the observed and unobserved (lost) signal components (see Sec. 5).
@@
  • The number of peaks ($N_{p}$) in the waveform . As explained before, for each peak, the radial position is reconstructed based on its drift time, and its $Z$ position is determined from the ratio between the left and right waveforms.
@@
Although the present simulation did not account for either the lifetime or the diffusion of the primaries, the most evident effects induced by these contributions on the simulated waveforms were examined. The reduction in the number of primaries during their migration was modelled according to the relation:
$N(t)=N(0)e^{-\frac{t}{\tau}},$ (3)
where $\tau$ is the lifetime of the electrons, $N(0)$ represents the number of primaries created at time $t$ = 0 prior to their drift, and $N(t)$ is the number of primaries remaining after a drift time $t$. For each hit, the associated charge loss is accounted for by multiplying the truncated form of this function with the corresponding elementary current.
As a byproduct of the GARFIELD++ results presented in Fig. 2, the spatial dispersion of the primaries (represented as the standard deviation of the distribution under the Gaussian hypothesis) was also extracted as a function of the distance travelled. This spatial dispersion was then converted into a temporal dispersion, $\sigma_{d}$. The latter, considered as an overall diffusion effect, was found to increase almost linearly with distance, starting from $\sigma_{d}=0~{}\mu$s for primaries generated near the anode, up to $\sigma_{d}\approx 4.2~{}\mu$s for those generated near the cathode. The complete waveform, incorporating diffusion effects, was obtained by convolving each elementary current function with a Gaussian whose standard deviation corresponds to the starting point of the hit.
For illustration, the deformation of the waveform shown in Fig. 3 is presented in Fig. 8, with parameters $\tau=10~{}\mathrm{ms}$ and $\sigma_{d}=2.7~{}\mu\mathrm{s}$ (at $R=26.5~{}\mathrm{cm}$). Further studies as a function of the interaction distance suggest that attenuation is the most significant detrimental effect. However, the signal can still be effectively decomposed as long as $\tau$ remains greater than the maximum drift time. These preliminary observations motivated the initiative not to include these two effects at this stage of the study.
Figure 8: Signal deformation under the combined effects of attenuation and diffusion, assuming $\tau=10~{}\mathrm{ms}$ and $\sigma_{d}=2.7~{}\mu\mathrm{s}$ (at $R=26.5~{}\mathrm{cm}$).
@@ 5 Energy reconstruction
  • The synthesis method also allows calculation of the reconstructed energy deficit. For each cluster of an event, identified by the decomposition algorithm at a given radial position $R$ (i.e. at a given drift time), the energy deficit is determined by multiplying the value of the lost charge function at that position by the number of primaries estimated at the same point. Summing over the different clusters enables the correction of the energy deficit due to the varying initial radial positions of the primaries in the event (see Fig. 10). To demonstrate the effectiveness of the method, Fig. 11 presents a comparison between the total reconstructed charge, $Q_{r}$, and the reconstructed charge after applying the decomposition algorithm, $Q_{synthesized}$, without calibration. The comparison is shown as a function of the mean radial position of the signal.
Figure 10: Instantaneous current (blue) induced by an elementary charge and corresponding lost charge (orange) as a function of the drift time. For illustration, the marker indicates the energy loss associated with a primary created at $R$ = 35.38 cm, and the corresponding waveform is shown by the blue area.
Figure 11: Reconstructed charge $Q_{r}$ (without the lost charge) and $Q_{synthesized}$ (including the lost charge) as a function of the signal mean radial position $R$ when no calibration procedure is applied (simulated $\beta\beta 0\nu$ events).
Figure 12: Reconstructed charge $Q_{synthesized}$ from the synthesis method as a function of the mean radial position of the signal $R$ before and after the calibration procedure explained in the text (simulated $\beta\beta 0\nu$ events).
@@
The study highlights the critical importance of experimental energy calibration of the detector due to the dependence of the signal on the radial position $R$. This calibration must also be implemented in real operation, and two potential solutions are currently under investigation. The first option involves using a dedicated point-like alpha source deployed at specific radial positions, though this could potentially disturb the drift electric field. Another attractive solution is to exploit the diffuse radon dissolved in the gas.
The energies reconstructed before and after the calibration procedures previously discussed are shown in Fig. 13. A Gaussian fit of the distribution obtained after energy correction shows a resolution of 0.1% FWHM for both methods (intrinsic resolution of the method since stochastic fluctuation on the number of created charges is not included at this stage). Although both approaches yield similar results, the synthesis routine is deemed more satisfactory from a methodological perspective. Consequently, this method was used for energy estimation in this article.
@@
To perform a realistic sensitivity study, the energy resolution must account for additional uncertainties. At this stage, effects due to fluctuations in the number of electrons or charge loss from attachment to gas impurities during drift have not been included. A separate statistical error of 0.3% FWHM is expected, assuming a Fano factor of approximately 0.15 Nygren:2011zz . Additionally, for real data, the influence of front-end electronics (amplifier response function) or digital processing (time shift of the sampling comb between the two ADCs) can also affect the energy resolution. It should be noted that these effects impact only the reconstructed energy and not the efficiency of the peak-finding algorithm or the selection variables derived from it. This was verified through a dedicated simulation of waveforms, which included a stochastic fluctuation in the number of electrons produced for each energy deposit.
@@
Figure 14: Selection cuts, indicated by the vertical red dashed line, on several observables, namely: number of reconstructed peaks 14(a), $\Delta R$ 14(b), $\Delta Z$ 14(c), $\Delta R\times\Delta Z$ 14(d), and larger radial energy deposit position 14(e). The distribution of such observables (after ROI selection) for $\beta\beta 0\nu$ (blue), Thorium 232 chain (orange) and Uranium 238 chain (green) are shown.
@@ 8.2 Composite material radioactivity
A specific composite material selection process is currently underway in collaboration with the IRT Jules Verne company. Various carbon fibres are being evaluated based on their radioactivity and mechanical strength to withstand pressure, alongside different epoxy glues to bind them. As mentioned in Sec. 2, a structure with a 1.5 cm thick carbon fibre-based composite material has been assumed. The decay chains of 238U and 232Th, were simulated uniformly within the carbon fibre layer. For the sensitivity studies, it was assumed that material radioactivity at the level of 10 $\mu$Bq/kg can be achieved. While this remains a critical point yet to be fully demonstrated, promising results have been obtained through the partnership with IRT Jules Verne. In any case, the presented background event rates per year can be scaled linearly with respect to different material radioactivity levels.
In the GEANT4 simulation, a conservative time window of 10 ms was assumed to separate events in the decay chain, considering a maximal drift time of approximately 5 ms. Given the half-lives of 164 $\mu$s for 214Po (from the 238U chain) and 300 ns for 212Po (from the 232Th chain), Bi-Po decays are treated as a single event in the detector simulation. For each decay chain, a sample of 10 million events was simulated. Only events depositing more than 2.2 MeV in the xenon active volume at the GEANT4 level were further processed. Based on the assumed activity of 10 $\mu$Bq/kg and the composite mass of approximately 330 kg, the generated statistics correspond to about 10 years of data for 232Th, and 7 years for 238U.
The total number of expected events per year in the ROI, before any selection cuts, can be computed, as well as the irreducible background events surviving the selection cuts described in Sec. 6. A summary is presented in Tab. 2, which shows that the total number of background events from the composite vessel’s radioactivity is approximately 1.2 per year. All the selected events in the 232Th decay chain originate from 208Tl decays, while the events from the 238U decay chain are from 214Bi decays. In both cases, a $\gamma$ deposits all the energy in a confined region of the detector, with no multi-Compton topology to enable event rejection.
@@
8.3 Anode radioactivity
The same simulation described in Sec. 8.2 was used to estimate the background contribution from the anode activity. As explained in Sec. 2, the use of a hollow tube with a thickness of 1 mm minimises the anode mass to less than 200 g. A statistical sample of 1 million events was required to observe any events passing the selection cuts, corresponding to approximately $1\times 10^{-3}$ events per year, under the assumption of an activity of 10 $\mu$Bq/kg, as in Sec. 8.2. Consequently, the background associated with the anode can be considered negligible, even if the estimated activity is increased by a factor of 100, i.e., assuming an activity of 1 mBq/kg.
8.4 Lead radioactivity
The lead shielding used to screen the detector from external gamma radiation could also introduce background due to contamination from $\mathrm{{}^{238}U}$, $\mathrm{{}^{232}Th}$, and their decay chains. The impact of this background was evaluated, assuming activities of $12~{}\mu\mathrm{Bq/kg}$ for uranium and $4~{}\mu\mathrm{Bq/kg}$ for thorium Leonard:2007uv . It should be noted that these values represent upper limits constrained by measurement sensitivity, and the actual contamination levels may be lower.
Through the full analysis chain, this background was found to contribute a non-negligible $3.5$ events per year for uranium and $1.9$ events per year for thorium. These contributions arise primarily from $\mathrm{{}^{214}Bi}$ and $\mathrm{{}^{208}Tl}$, produced within the first few inner centimetres of the lead shielding. However, if the first $\sim 5~{}\mathrm{cm}$ of lead are excluded, the background is significantly reduced to an almost negligible level of $0.1$ events per year. Again, this represents an upper limit.
To mitigate the contribution of the shielding, several options are being considered for the final setup. One option is to include a copper layer a few centimetres thick. Ultra-low-radioactivity copper, with contamination levels below $1~{}\mu\mathrm{Bq/kg}$, is commercially available from suppliers such as Aurubis or Mitsubishi MAJORANA:2016lsk . Alternatively, a water-based passive veto could be employed.
In the present study, a contribution of $0.1$ events per year is assumed, with the most practical option to achieve this background level still under evaluation.